At a certain temperature, 0.980 mol SO3 is placed in a 4.00 L container.

2SO3(g)−⇀↽−2SO2(g)+O2(g)

At equilibrium, 0.180 mol O2 is present. Calculate Kc.
Answer #1

Answer:

Step 1: Explanation

Kc is defined as the amounts of products and reactants present at equilibrium in a reversible chemical reaction at a given temperature

Step 2: Write the balanced the chemical equation

2SO3 (g) <--------------> 2SO2 (g)+O2(g)

Step 3: Calculate the Concentrations by dividing the given moles by volume of container.

Given,

Initial SO3 = moles / volume = (0.980 moles / 4 litre ) = 0.245 M

as well the concentration of O2 at equilibrium = ( 0.180 moles / 4 L ) = 0.045 M

Step 3: Write the ICE table

2SO3 (g) <--------------> 2SO2 (g) + O2(g)

Initial amount 0.245 0 0

Change amount    -2x +2x +x

Equilibrium amount 0.245-2x 2x x

Since the value of O2 at equlibrium is given [O2] = 0.045 M = [x]

hence value of x = 0.045 M

Hence all value at equilibrium is

[SO3] = 0.245 -2x = 0.245 -(2 × 0.045 ) =0.155

[SO2] = 2x = 2 × 0.045 = 0.09

[O2] = x = 0.045

Step 4: Calculate the Kc

2SO3 (g) <--------------> 2SO2 (g) + O2(g)

The Kc expression will be

Kc = [SO2]2 × [O2] / [SO3]2

on substituting the value

Kc= ( 0.09)2×(0.045) / (0.155)2 = 0.0151717 ≈  1.52 ×10-2

Hence, the value of Kc = 0.0151717 ≈  1.52 ×10-2